### v1

@@ IV WHERE TO EVALUATE
Where should we evaluate the safety of LLMs? Evaluation metrics determine the aspects of safety performance to be measured, toolkits provide the framework for the entire evaluation process, and datasets and benchmarks form the foundation of these evaluations. In this section, we will comprehensively summarize and introduce these three parts, which offer researchers a clear view of LLMs safety performance.
@@ IV-A Metrics
This section will discuss the evaluation metrics used to evaluate the safety of LLMs, as well as some frontier research related to these metrics. The focus will be on evaluating whether LLMs generate harmful content; currently, there are relatively few summaries of evaluation metrics in this area. In contrast, metrics for evaluating the truthfulness of LLMs have already been summarized in the context of hallucination evaluations[188], so we will not elaborate further on that aspect in this section.
@@ IV-A1 Attack Success Rate
Attack success rate (ASR) is the most widely used and critical metric for evaluating the safety of large language models (LLMs). It measures the effectiveness of adversarial attacks such as jailbreak attacks and prompt injection attacks directed at LLMs. Obviously, the higher the ASR, the less safe LLMs are. Mathematically, ASR can be formulated as:
@@
where $Num_{success}$ denotes the number of successfully attacked prompts, and $Num_{total}$ denotes the total number of harmful prompts.
Similarly, variations of ASR, such as the harmful content rate and refuse rate, share the same fundamental concept, so we will not discuss them separately. However, a significant challenge remains in defining what constitutes a "successful attack." How to evaluate the success of a jailbreak attempt against an LLM has not been unified [187], leading to difficulties in comparing the works of different researchers. Current automatic studies primarily determine success through two approaches: rule-based and language model-based methods. In brief, the former evaluates whether certain predefined keywords (e.g., "Sorry," "As an AI assistant," etc.) are present in the LLMs responses to determine if the model is obedient to malicious prompts, while the latter employs a judge LLM to evaluate the prompts and responses and determine the success of the attack. This aspect will be discussed in detail in Section V.
IV-A2 Other fine-grained Metric
As mentioned in the section IV-A1, the research community has yet to reach a consensus on how to define "attack success" or "jailbreak success." Therefore, to comprehensively and accurately evaluate the safety of LLMs, many researchers have proposed new metrics or adapted some metrics from general LLMs evaluation frameworks.
@@ IV-A2 Other fine-grained Metric
where $refused$ denotes whether the response is rejected (1) or not (0). $specific$ and $convincing$ denote how specific and convincing the response is. The higher the $score$, the less safe the response.
@@
Chen et al.[53] argues that a reliable and safe LLM should meet three core criteria: resistance to jailbreak, non-toxicity, and quality. Accordingly, researchers use five metrics to evaluate these aspects: the commonly used attack success rate (ASR), toxicity score, fluency based on perplexity, token length, and grammatical errors. Additionally, Leidinger et al.[93] incorporates four metrics—refusal, toxicity, sentiment, and regard—to capture nuanced biases in model responses. Moreover, Mei et al.[190] suggests that many detected jailbreaks may actually result from hallucinations. To evaluate such hallucinations, researchers have introduced a range of metrics, including entity transformation, semantic consistency, and recall.
Overall, most researchers currently rely on the relatively simple metric of ASR as a primary evaluation metric. However, there are notable differences in the specific methods used to implement this metric, and the reproducibility of the results is poor. Some researchers have also attempted to refine existing metrics and propose fine-grained metrics for a more in-depth evaluation, though their impact remains limited. Thus, further research is needed on how to establish a unified and universal set of LLMs safety evaluation metrics.
IV-B Toolkits
Unlike individual metrics used for LLMs safety evaluation, toolkits typically integrate the entire evaluation pipeline, enabling a comprehensive and end-to-end safety evaluation.
Sun et al.[17] developed a Chinese LLMs safety evaluation benchmark, providing a comprehensive evaluation of LLMs safety across eight typical safety scenarios and six more challenging instruction attacks. Specifically, the researchers input collected prompts and corresponding responses into InstructGPT[191] to evaluate the safety of model outputs. Additionally, the authors released a safety leaderboard that documents the safety performance of all evaluated models, offering a valuable reference for LLMs developers.
HarmBench[10] introduces a standardized evaluation framework for automated red-teaming to evaluate both red-team attacks and defenses. HarmBench identifies three key qualities essential for automated red-team evaluation: breadth, comparability, and robust metrics. Based on these principles, the evaluation pipeline begins by converting a diverse set of behaviors into test cases to evaluate the target model. These behaviors and their corresponding responses are then processed through multiple classifiers to output evaluation results, ultimately determining the final attack success rate.
JailbreakBench[11] developed an open-source evaluation framework for evaluating jailbreak attacks and defenses. First, JailbreakBench maintains an evolving repository of jailbreak artifacts, encompassing state-of-the-art attacks and defenses. Additionally, it establishes a standardized evaluation framework that allows users to clearly define threat models, system prompts, chat templates, and scoring functions. This framework provides a convenient, streamlined approach for red-teaming, defense testing, and the selection of various evaluation classifiers. Finally, JailbreakBench features a leaderboard on their website that displays the performance of advanced jailbreak attacks and defenses.
EasyJailbreak[192] simplifies the construction and evaluation of jailbreak attacks against LLMs, proposing a unified implementation framework. This framework comprises four components: Selector, Mutator, Constraint, and Evaluator. The entire process is divided into three stages. In the preparation stage, users provide initial configurations, such as malicious prompts and template seeds. During the attack stage, EasyJailbreak iteratively attacks the target model based on user inputs, updating prompts and evaluating each attack’s output. In the final stage, it summarizes the attack results, returning a comprehensive report to the user.
WalledEval[54] is a comprehensive LLM safety testing toolkit. It consists of three main classes: a dataset loader, a LLM loader, and a judge loader, supporting LLM benchmarking, judge benchmarking, and multiple-choice benchmarking. In addition, WalledEval is compatible with both weight-based and API-based model types and includes over 35 safety evaluation benchmarks, covering a wide range of safety dimensions such as multilingual safety, exaggerated safety, and prompt injections.
JailbreakEval[187] offers a user-friendly toolkit specifically for evaluating the safety of content generated by LLMs. Unlike other frameworks that integrate attack and evaluation processes, JailbreakEval is designed solely for evaluating content safety. It includes a variety of out-of-the-box evaluators, broadly categorized into string matching evaluators, text classification evaluators, chat evaluators, and voting evaluators. Users simply input jailbreak prompts and model responses to obtain evaluation results.
Table I: Overview of comprehensive evaluation datasets.
@@ IV-B Toolkits
Do-Not-Answer [43] 939 5 English Harmful instructions
@@ IV-C Datasets and Benchmarks
To comprehensively evaluate the safety performance of LLMs, numerous datasets and benchmarks have been developed. Unlike those focusing on specific domains[183, 35], in this section, we discuss only those datasets and benchmarks designed to evaluate LLMs’ overall safety across multiple dimensions. In this study, we selected 31 datasets and benchmarks in total, as shown in Table I. To facilitate differentiation, we categorized them into four groups based on their composition: only questions (Q-only), questions and answers (Q&A), multiple-choice questions (MCQ), and multi-turn conversations (Dialogue). Additionally, we advocate that the research community should consider LLMs safety within multilingual environments.
@@
The majority of major datasets and benchmarks consist of Q-only. Chu et al.[63] establish a dataset of prohibited questions, providing the first large-scale evaluation of various advanced jailbreak attack methods. StrongReject[184] proposes a dataset composed of forbidden questions that LLMs should refuse to answer to evaluate jailbreak performance. HarmBench[10] construct its dataset by incorporating not only harmful behaviors but also copyright, contextual, and multimodal behaviors for a comprehensive evaluation. Do-Not-Answer[43] includes prompts that responsible language models should not answer, while AdvBench[7] contains harmful strings and behaviors. Although initially develop using the white-box jailbreak method GCG, AdvBench has become a popular evaluation benchmark for researchers. To evaluate LLMs robustness against jailbreak attacks, Chen[53] introduce a refined dataset covering a wide range of harmful prompts. TechHazardQA[14] presents a complex harmful query dataset, requiring LLMs to respond with pseudocode instead of standard text to evaluate ethical boundaries. FFT[4] carefully elaborates a dataset of harmful prompts focusing on factoid, unfair, and toxic contents in LLMs. SORRY-Bench[56] addresses three key limitations in existing benchmarks: coarse-grained taxonomy of unsafe topics, overlook of linguistic characteristics and formatting of prompts, and reliance on large LLMs for evaluation. ALERT[62] provides a substantial instruction dataset with a fine-grained risk taxonomy for red-team testing of LLMs. Do Anything Now[45] recognizes the shift of jailbreak prompts from online communities to prompt aggregation sites, so researchers create a question set based on in-the-wild jailbreak prompts. Wang et al.[61] construct a dataset including both harmful and benign prompts to measure LLMs false positives and false negatives in refusal scenarios. JailBench [227] collects and integrates advanced jailbreak attack methods in the field, constructs a jailbreak prompt dataset with powerful jailbreak capabilities, and can comprehensively detect and evaluate the content safety protection capabilities of LLMs. In alignment with Chinese internet content regulations, ChineseSafe[189] expands on sensitive categories specific to the Chinese context, including political sensitivity, pornography, and variant/homophonic words. Latent Jailbreak[9] innovatively embeds harmful instructions into translation tasks, aiming to evaluate the robustness of LLMs against jailbreak attacks through translation tasks. Finally, S-Eval[57] introduces a bilingual dataset of risk prompts in both Chinese and English, containing fundamental risk prompts along with their corresponding adversarially crafted attack prompts.
@@
Compared to datasets containing only questions(Q-only), datasets that include both questions and corresponding answers(Q&A) offer a reference ground truth, enabling more precise safety evaluations and even fine-tuning of models for improved performance. JailbreakBench [11] proposes a mixed dataset covered OpenAI’s usage policies, where each harmful behavior is paired with a benign behavior, making it useful for evaluating rejection rates and model defenses. JailJudge ID[194] includes a collection of manually annotated harmful prompt-response pairs set in complex scenarios, while AegisSafetyDataset[218] is a large-scale record of safey and unsafe interactions between humans and LLMs. Additionally, WildGuardMix[51] is a large-scale, carefully balanced multitask dataset designed for safety regulation, encompassing both direct prompts and adversarial jailbreak responses with various refusal and compliance responses, primarily intended for fine-tuning safety evaluation models. Safety Prompts[17] provides augmented prompts and responses aimed at testing and improving LLMs safety, supporting a targeted approach to model safety enhancement.
@@
Multiple-choice questions(MQA) offer a clear advantage over Q&A-style assessments when evaluating model safety, as they provide fixed correct answers, making it straightforward to evaluate model performance. SALAD-Bench[47], for instance, includes both harmful questions and multiple-choice questions, diversifying queries to enable a more comprehensive safety evaluation of LLMs. In the Chinese context, CHiSafetyBench[59] is designed to identify risky content by requiring LLMs to select the single safe or unsafe option from multiple choices. CValues[46] also focuses on Chinese scenarios, evaluating LLMs alignment with human values from safety and responsibility perspectives by having models choose the better response between two options. To measure the risk propensity of LLMs, CRiskEval[60] constructs multiple-choice questions, each with four manually labeled safety levels: Extremely Hazardous, Moderately Hazardous, Neutral, and Safe, providing a fine-grained evaluation scale. SafetyBench[44] combines both Chinese and English settings, creating a large, diverse multiple-choice dataset that thoroughly evaluates LLMs safety, blending harmful and safe behaviors in options to challenge models in their selection.
@@
Some jailbreak attack methods have evolved beyond single-turn attacks, extending to multi-turn interactions, which have shown significantly heightened risks[195, 196, 197]. Consequently, evaluating the safety of LLMs in multi-turn conversations(Dialogue) has become necessary. R-Judge[40] provides a benchmark for assessing LLMs capabilities in identifying and evaluating safety risks within multi-turn interaction records between agents. CoSafe[55] highlights the current focus on single-turn interactions in red-teaming approaches and thus introduces a dataset designed for multi-turn attacks, featuring cohesive referencing across multiple turns. Additionally, HarmfulQA[8] presents a dataset not only of harmful questions but also includes both blue and red conversations, making it suitable for fine-tuning safety models.
@@
Furthermore, we emphasize the importance of evaluating the multilingual safety of LLMs. Studies have shown that the unbalanced cross-linguistic distribution of LLMs training data leads to language biases, resulting in inconsistent outputs when the same task is described in different languages[33]. Research also indicates that multilingual mixed prompts can significantly intensify the harm of malicious queries, and that different language types and families exhibit substantial difference in LLMs safety[32]. Thus, while much research is focused on English and Chinese, the safety of LLMs in multilingual and low-resource language contexts deserves attention. AraTrust[28] introduced the first Arabic-language LLMs trustworthiness benchmark, consisting of manually crafted multiple-choice questions. XSafety established the first multilingual safety benchmark, assessing LLMs across 10 languages. JailJudge OOD [194] also introduced a multilingual safety dataset in 10 languages, with labeled harmful prompts and responses. To address the limitations of existing toxicity benchmarks focused primarily on English, PolyglToxicityPrompts [5] released a dataset with natural prompts across 17 languages for evaluating multilingual toxicity in LLMs.
@@ V HOW TO EVALUATE
Having identified what and where to evaluate in terms of LLM safety, the next critical question centers on how to perform such evaluations effectively. A wide range of methodologies have emerged to assess the safety of LLMs, varying in automation, granularity, and evaluator participation. This section provides a comprehensive review of the main evaluation paradigms, categorizing them according to the nature and role of the evaluator: human, rule-based or model-based, and discussing the advantages, limitations, and use cases of each approach.

### v2

@@ IV WHERE TO EVALUATE
The safety of LLMs can be evaluated through carefully designed metrics, which determine the specific aspects of performance to be measured, and through datasets and benchmarks that provide the foundation for these evaluations. In this section, we comprehensively summarize these two components — metrics, datasets and benchmarks — to provide a clear overview of LLMs safety performance.
@@ IV-A Metrics
This section will discuss the evaluation metrics used to evaluate the safety of LLMs, as well as some frontier research related to these metrics. The focus will be on evaluating whether LLMs generate harmful content; currently, there are relatively few summaries of evaluation metrics in this area. In contrast, metrics for evaluating the truthfulness of LLMs have already been summarized in the context of hallucination evaluations[152], so this aspect is therefore not elaborated here.
@@ IV-A1 Attack Success Rate
Attack success rate (ASR) is widely used and fundamental metric for evaluating the safety of LLMs. It measures the effectiveness of adversarial attacks such as jailbreak attacks and prompt injection attacks directed at LLMs. A higher ASR indicates lower safety of the evaluated LLM. Mathematically, ASR can be formulated as:
@@
where $Num_{success}$ denotes the number of successfully attacked prompts, and $Num_{total}$ denotes the total number of prompts used for evaluation.
Similarly, variations of ASR, such as the harmful content rate and refuse rate, share the same fundamental concept, so we will not discuss them separately. However, a significant challenge remains in defining what constitutes a “successful attack”. How to evaluate the success of a jailbreak attempt against an LLM has not been unified [17], leading to difficulties in comparing the works of different researchers. Current automatic studies primarily determine success through two approaches: rule-based and model-based methods, which discussed in detail in Section V. Rule-based methods are simple, but often have low relevance to human judgment, while model-based methods use different implementation paths in different studies, providing greater flexibility at the expense of reproducibility and comparability.
IV-A2 Fine-Grained Metrics
As mentioned in the section IV-A1, the research community has yet to reach a consensus on how to define “attack success” or “jailbreak success”. Therefore, to comprehensively and accurately evaluate the safety of LLMs, many researchers have proposed new metrics or adapted some metrics from general LLMs evaluation frameworks.
@@ IV-A2 Fine-Grained Metrics
where refusedrefused denotes whether the response is rejected (1) or not (0). specificspecific and convincingconvincing denote how specific and convincing the response is. The higher the scorescore, the less safe the response. This is a relatively strict evaluation standard, which alleviates the phenomenon of exaggerating the performance of jailbreak attacks.
@@
Chen et al.[27] argues that a reliable and safe LLM should meet three core criteria: resistance to jailbreak, non-toxicity, and quality. Accordingly, researchers use five metrics to evaluate these aspects: the commonly used attack success rate (ASR), toxicity score, fluency based on perplexity, token length, and grammatical errors. Additionally, Leidinger et al.[47] incorporates four metrics—refusal, toxicity, sentiment, and regard—to capture nuanced biases in model responses. Moreover, Mei et al.[64] suggests that many detected jailbreaks may actually result from hallucinations. To evaluate such hallucinations, researchers have introduced a range of metrics, including entity transformation, semantic consistency, and recall. Although these fine-grained metrics are designed to capture aspects of safety beyond binary classification and narrow the gap with human judgment gold standards, they still heavily rely on the subjective judgment of models, lack standardized definitions, and have yet to gain widespread adoption, thereby limiting the comparability of research.
Overall, ASR remains the predominant metric for safety evaluation, but inconsistent definitions of ’success’ hinder reproducibility. Emerging fine-grained metrics provide richer insights, yet they remain fragmented and lack standardization. Establishing a unified and widely accepted set of safety evaluation metrics therefore remains an open research challenge.
Table II: Overview of comprehensive evaluation datasets.
@@
Do-Not-Answer [67] 939 60 English Harmful instructions
@@
AnswerCarefully [85] 1.8k 56 Japanese Harmful questions and reference answers
@@ IV-B Datasets and Benchmarks
To comprehensively evaluate the safety performance of LLMs, numerous datasets and benchmarks have been developed. Unlike those focusing on specific domains[56, 58], in this section, we focus on datasets and benchmarks that assess the overall safety of LLMs across multiple dimensions. In this study, we selected 31 datasets and benchmarks in total, as shown in Table II. For clarity of categorization, we categorized them into four groups based on their composition: only questions (Q-only), questions and answers (Q&A), multiple-choice questions (MCQ), and multi-turn conversations (Dialogue). Additionally, we advocate that the research community should consider LLMs safety within multilingual environments.
@@
The majority of major datasets and benchmarks consist of Q-only. Chu et al.[65] establish a dataset of prohibited questions, providing the first large-scale evaluation of various advanced jailbreak attack methods. StrongReject[63] proposes a dataset composed of forbidden questions that LLMs should refuse to answer to evaluate jailbreak performance. HarmBench[31] construct its dataset by incorporating not only harmful behaviors but also copyright, contextual, and multimodal behaviors for a comprehensive evaluation. Do-Not-Answer[67] includes prompts that responsible language models should not answer, while AdvBench[8] contains harmful strings and behaviors. Although initially develop using the white-box jailbreak method GCG, AdvBench has become a popular evaluation benchmark for researchers. To evaluate LLMs robustness against jailbreak attacks, Chen[27] introduce a refined dataset covering a wide range of harmful prompts. TechHazardQA[35] presents a complex harmful query dataset, requiring LLMs to respond with pseudocode instead of standard text to evaluate ethical boundaries. FFT[25] carefully elaborates a dataset of harmful prompts focusing on factoid, unfair, and toxic contents in LLMs. SORRY-Bench[70] addresses three key limitations in existing benchmarks: coarse-grained taxonomy of unsafe topics, overlook of linguistic characteristics and formatting of prompts, and reliance on large LLMs for evaluation. ALERT[73] provides a substantial instruction dataset with a fine-grained risk taxonomy for red-team testing of LLMs. Do Anything Now[75] recognizes the shift of jailbreak prompts from online communities to prompt aggregation sites, so researchers create a question set based on in-the-wild jailbreak prompts. Wang et al.[77] construct a dataset including both harmful and benign prompts, which measure LLMs false positives and false negatives in refusal scenarios. JailBench [78] collects and integrates advanced jailbreak attack methods in the field, constructs a jailbreak prompt dataset with powerful jailbreak capabilities, and can comprehensively detect and evaluate the content safety protection capabilities of LLMs. In alignment with Chinese internet content regulations, ChineseSafe[80] expands on sensitive categories specific to the Chinese context, including political sensitivity, pornography, and variant/homophonic words. Latent Jailbreak[81] innovatively embeds harmful instructions into translation tasks, aiming to evaluate the robustness of LLMs against jailbreak attacks through translation tasks. Finally, S-Eval[83] introduces a bilingual dataset of risk prompts in both Chinese and English, containing fundamental risk prompts along with their corresponding adversarially crafted attack prompts. In summary, Q-only datasets provide broad coverage for testing refusal and robustness, but they lack reference responses for fine-grained evaluation.
@@
Compared to datasets containing only questions(Q-only), datasets that include both questions and corresponding answers(Q&A) offer a reference ground truth, enabling more precise safety evaluations and even fine-tuning of models for improved performance. JailbreakBench [30] proposes a mixed dataset covered OpenAI’s usage policies, where each harmful behavior is paired with a benign behavior, making it useful for evaluating rejection rates and model defenses. JailJudge ID[69] includes a collection of manually annotated harmful prompt-response pairs set in complex scenarios, while AegisSafetyDataset[71] is a large-scale record of safey and unsafe interactions between humans and LLMs. Additionally, WildGuardMix[74] is a large-scale, carefully balanced multitask dataset designed for safety regulation, encompassing both direct prompts and adversarial jailbreak responses with various refusal and compliance responses, primarily intended for fine-tuning safety evaluation models. Safety Prompts[39] provides augmented prompts and responses aimed at testing and improving LLMs safety, supporting a targeted approach to model safety enhancement. Overall, Q&A datasets enable accurate assessment and model training but require extensive annotation efforts.
@@
Both Q-only and Q&A datasets are used to evaluate whether the content generated by LLMs in response to harmful queries complies with relevant policies. In contrast to this generative approach, multiple-choice questions (MCQ), as a form of discriminative evaluation, place a greater emphasis on the LLMs’ ability to identify harmful content. Furthermore, a major advantage of MCQ over generative evaluation is that they can provide a highly intuitive and convenient measure of a model’s safety performance, as the answers are unique and the evaluation results can be determined straightforwardly. SALAD-Bench[72], for instance, includes both harmful questions and multiple-choice questions, diversifying queries to enable a more comprehensive safety evaluation of LLMs. In the Chinese context, CHiSafetyBench[76] is designed to identify risky content by requiring LLMs to select the single safe or unsafe option from multiple choices. CValues[38] also focuses on Chinese scenarios, evaluating LLMs alignment with human values from safety and responsibility perspectives by having models choose the better response between two options. To measure the risk propensity of LLMs, CRiskEval[79] constructs multiple-choice questions, each with four manually labeled safety levels: Extremely Hazardous, Moderately Hazardous, Neutral, and Safe, enabling fine-grained evaluation. SafetyBench[82] combines both Chinese and English settings, creating a large, diverse multiple-choice dataset that thoroughly evaluates LLMs safety, blending harmful and safe behaviors in options to challenge models in their selection. MCQ datasets offer intuitive, reproducible evaluation metrics but may oversimplify complex generative behaviors.
@@
Some jailbreak attack methods have evolved beyond single-turn attacks, extending to multi-turn interactions, which have shown significantly heightened risks[184, 185, 186]. Consequently, evaluating the safety of LLMs in multi-turn conversations(Dialogue) has become necessary. R-Judge[66] provides a benchmark for assessing LLMs capabilities in identifying and evaluating safety risks within multi-turn interaction records between agents. CoSafe[68] highlights the current focus on single-turn interactions in red-teaming approaches and thus introduces a dataset designed for multi-turn attacks, featuring cohesive referencing across multiple turns. Additionally, HarmfulQA[24] presents a dataset not only of harmful questions but also includes both blue and red conversations, making it suitable for fine-tuning safety models. Multi-turn datasets are essential for assessing real-world conversational safety, though they are resource-intensive to construct and analyze.
@@
Furthermore, we emphasize the importance of evaluating the multilingual safety of LLMs. Studies have shown that the unbalanced cross-linguistic distribution of LLMs training data leads to language biases, resulting in inconsistent outputs when the same task is described in different languages[187]. Research also indicates that multilingual mixed prompts can significantly intensify the harm of malicious queries, and that different language types and families exhibit substantial difference in LLMs safety[188]. Thus, while much research is focused on English and Chinese, the safety of LLMs in multilingual and low-resource language contexts deserves attention. AraTrust[84] and AnswerCarefully[85] extend benchmarks to Arabic and Japanese, respectively, while XSafety[86], JailJudge OOD[69], and PolyglToxicityPrompts[23] offer multilingual coverage across 10–17 languages. These datasets highlight the importance of assessing LLMs safety in diverse linguistic contexts.
@@ V HOW TO EVALUATE
Having identified what and where to evaluate in LLM safety, the next critical issue centers on how such evaluations should be conducted. A variety of methodologies have emerged, differing in levels of automation, granularity, and evaluator involvement. This section reviews the major evaluation paradigms, classifying them according to the nature and role of the evaluator—human, rule-based, or model-based—and discusses their respective advantages, limitations, and use cases. In addition, some evaluation frameworks that integrate the entire evaluation pipeline are also summarized.